OpenStudy (littleheron):

Solve for y. 4x – 2y = 10 y = 2x – 5 y = -4x – 5 y = 2x + 5 y = -2x – 5

OpenStudy (just_one_last_goodbye):

Need help?

OpenStudy (littleheron):

Yep

OpenStudy (ryuga):

get y by itself

OpenStudy (ryuga):

2 -2 --- 0 = x 10 -2 --- 8 x = 8

OpenStudy (anonymous):

Hey , guys can I help out?

OpenStudy (littleheron):

sure

OpenStudy (ryuga):

already go thelp

OpenStudy (ryuga):

*got

OpenStudy (anonymous):

Ok so we will have to solve for y..correct?

OpenStudy (just_one_last_goodbye):

No direct answers please :)

OpenStudy (just_one_last_goodbye):

@Ryuga your answer is wrong

OpenStudy (just_one_last_goodbye):

as that does not match any answer choices ^_^

OpenStudy (anonymous):

Mr. obvious

pooja195 (pooja195):

Basically when you are solving for y 4x – 2y = 10 We want y by itself subtract 4x from both sides -2y=4x+10 now divide everything by -2 to isolate the y

OpenStudy (littleheron):

xD

OpenStudy (anonymous):

thank you pooja

pooja195 (pooja195):

:) yw ^_^

OpenStudy (anonymous):

ok so wat do u guys vote 4

OpenStudy (anonymous):

a ,b , c , or d now if u r not really sure then don't answer

OpenStudy (littleheron):

i vote a

OpenStudy (anonymous):

i vote a too...ok 2 votes 4 a any rejects

OpenStudy (anonymous):

ok no rejects so it is a go ahead